OpenStudy (anonymous):

another integration question. integrate (9x^2)/(sqrt (1-x^2)) dx

sam (.sam.):

I think you need trig sub

OpenStudy (anonymous):

i've try but don't get it :(

sam (.sam.):

try substitute x=sin u and x=cos u du Then \[\sqrt{1-x^2}=\sqrt{1-\sin ^2u }=\cos u\] \[u=\sin ^{-1}(x)\] \[9\int\limits \sin ^2(u) \, du\] Use half angle formula \[9\int\limits \left(\frac{1}{2}-\frac{1}{2} \cos (2 u)\right) \, du\] \[\text{}=9\int\limits \frac{1}{2} \, du-\frac{9}{2}\int\limits \cos (2 u) \, du\] for cos2u, use v=2u dv=2du \[9\int\limits \frac{1}{2} \, du-\frac{9}{4}\int\limits \cos v \, dv\] integrate separately, \[\frac{9 u}{2}-\frac{9 \sin (s)}{4}+c\] sub back those leftovers \[\frac{9}{2} \sin ^{-1}(x)-\frac{9}{2} x \sqrt{1-x^2}\]

OpenStudy (anonymous):

if we don't use trigo can get the answer or not?

sam (.sam.):

i don't think so

OpenStudy (anonymous):

but the answer is -6(1-x^2)^ (1/2) +c

sam (.sam.):

you sure you wrote the question correctly? if yes then im not sure, but this is my work

OpenStudy (anonymous):

\[\int\limits_{}^{} (9x^2)/\sqrt{1-x^2} dx\]